We wish to test the claim that μ > 32 at a level of significance of α = 0.05.
The sample statistics are n = 50, and x-bar = 32.3. Moreover, it is known that the
population mean is s = 12.


Compute the test statistic, rounded to two decimal places.

Group of answer choices

3.11

0.98

2.31

1.77
Answer #1

test statistic formula

\frac{(\bar{x} -\mu)}{\sigma/\sqrt{n}}

\frac{(32.3-32)}{12/\sqrt{50}}

0.1767

0.177

please like ??